OpenStudy (anonymous):

What is the standard form equation of the line passing through (–2, 8) with a slope of 2?

OpenStudy (anonymous):

The equation for this type of problem is: y-y1=m(x-x1) plugging in you get: y-8=2(x--2) which translates to y-8=2(x+2) and you can simplify more if need be.

OpenStudy (experimentx):

write the equation of circle in straight intercept form. put the vales of x and y from the point, determine the constant. replace constant from slope intercept form. you will have your equation.

Directrix (directrix):

From above, y-8=2(x+2) y - 8 = 2x + 4 --> using the Distributive Property y = 2x + 4 + 8 y = 2x + 12 Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. y = 2x + 12 -2x + 1y = 12 Multiply through by -1 2x - 1y = -12